OpenStudy (anonymous):

Exponent Questions... The exponents are the -2 and 2 3g-2b2 a. 3b2/g2 --- both 2 are exponents b. 3g2b-2 --- 2 and -2 are the exponents c. 3gb-4 --- negative four is the exponent d. b2/3g2 --- both 2 are exponents

OpenStudy (phi):

use ^ so 3g^(-2) b^2

OpenStudy (anonymous):

Thats much easier! Thank you:) Do you know the answer?

OpenStudy (phi):

remember this rule: if you have \[ a^{-b} = \frac{1}{a^b} \] flip the number and change the sign of the exponent also \[ \frac{1}{a^{-b}} = a^b\]

OpenStudy (anonymous):

I think its a or d...but im not sure which one...

OpenStudy (phi):

look at \[ 3g^{-2} b^2 \] look just at the g^(-2) can you use the rule to make the exponent positive ? (flip it!)

OpenStudy (anonymous):

So its b^2/3g^2?

OpenStudy (phi):

yes, except the 3 is not part of the g. the 3 stays up top

OpenStudy (phi):

do you remember order of operations PEDMAS ? parens, exponents then divide or multiply

OpenStudy (anonymous):

Ohhhhh so then its 3b^2/g^2

OpenStudy (phi):

yes

OpenStudy (anonymous):

Yay! Thank you so much for explaining that!

OpenStudy (anonymous):

Can you help me with another one?

OpenStudy (phi):

ok

OpenStudy (anonymous):

3/g^-2h^-3

OpenStudy (anonymous):

I have to simplify right?

OpenStudy (phi):

\[ \frac{3}{g^{-2} h^{-3} } \] ? I would flip and make the negative exponents positive

OpenStudy (phi):

first do the g and then the h

OpenStudy (anonymous):

To get